OpenStudy (anonymous):

the equation of the line tangent to the curve y=kx+8/k+x at x= -2 is y=x+4. what is the value of k?

OpenStudy (anonymous):

check your question , possibly a typo

OpenStudy (anonymous):

no, thats it word for word

OpenStudy (anonymous):

y= (k+1)x +(8/k) ( eqn of the "curve" , even though its a line )

OpenStudy (anonymous):

the only way to have a tangent to a line , is for the tangent to be the line

OpenStudy (anonymous):

so (k+1)x +(8/k) = x+4

OpenStudy (anonymous):

now we are told they are a tangent at x=-2, so we could sub x=-2 into the above to find k , but we could also sub any value of x we like , these lines have to be the same for all x

OpenStudy (anonymous):

but ill just sub x=-2 -2(k+1) + 8/k = 2 -2k(k+1) +8 = 2k -2k^2 -2k -2k +8 =0 -2k^2 -4k+8 =0 -2(k^2 +2k -4) =0

OpenStudy (anonymous):

k^2 +2k -4 =0

OpenStudy (anonymous):

how did you get that function?

OpenStudy (anonymous):

(k+1)^2 = 5 k +1 = +-sqrt(5) k = -1+-sqrt(5)

OpenStudy (anonymous):

not an answer

OpenStudy (anonymous):

well your teacher is stupid

OpenStudy (anonymous):

i doubt that you dont even know what the equation of the line tangent is

OpenStudy (anonymous):

this is a question from an ap test so there are no typos

OpenStudy (anonymous):

lol was the curve \[y=kx + \frac{8}{k+x}\] y =

OpenStudy (anonymous):

if it was................ :|

OpenStudy (anonymous):

use brackets when you are writting fractions on the net :|

OpenStudy (anonymous):

put the kx on top also

OpenStudy (anonymous):

sorry? didnt know

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

y=(kx+8)/k+x

OpenStudy (anonymous):

With the given info, one finds the tangent vector of the curve to be <1, -1> and the point of tangency to be (-2, 2). You can find k by plugging in (-2, 2).